OpenStudy (anonymous):

Which equation can be used to find the answer? Usman paints the entire length of a fence in a day. He paints the last 12 m of the 38 m fence in the afternoon. How many meters of the fence does he paint in the morning? A. f = 38 ÷ 2 f = 19 19 m B. f – 12 = 38 f = 50 50 m C. f = 12 · 2 f = 24 24 m D. f + 12 = 38 f = 26 26 m

OpenStudy (anonymous):

sub tract 12 from 38

RhondaSommer (rhondasommer):

actually in this equation you are going to have to add to equal 38.

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

yes ik u r solving for f

RhondaSommer (rhondasommer):

@usercode3rror take that down. you should know by now that no direct answers are allowed

OpenStudy (anonymous):

lol ik but my old accont is a auth

OpenStudy (anonymous):

Which equation can be used to find the answer? Dania marks off 18.7 m of a playing field for a sports festival. She needs to mark off a total of 29 m to finish the field. How many more meters does Dania have left to mark off? A. f = 29 ÷ 2 f = 14.5 14.5 m B. f – 18.7 = 29 f = 47.7 47.7 m C. f = 18.7 · 2 f = 37.4 37.4 m D. f + 18.7 = 29 f = 10.3 10.3 m

RhondaSommer (rhondasommer):

@sophissmart63which one do you think is correct

RhondaSommer (rhondasommer):

also; this belongs in another question. Thanks :)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@RHONDASOMMER the moderaters dont care if people do this it helps the site like this

RhondaSommer (rhondasommer):

how does B make sense? the finishing number is 47 M but you only have a 29 M field

OpenStudy (anonymous):

A

RhondaSommer (rhondasommer):

no actually you are wrong @usercode3rror . they do care and will take action if you do it too many times. Read the code of conduct it you cant get it together.

OpenStudy (anonymous):

@them n y do i c a auth doing it all the time

OpenStudy (anonymous):

@tkhunny

RhondaSommer (rhondasommer):

@sophissmart6330 why would you divide the field into 2? It is nowhere in the problem. instead of just guessing; please put logic into your answers

OpenStudy (anonymous):

@rhondasommer lol

RhondaSommer (rhondasommer):

why lol

OpenStudy (anonymous):

lol i know a few people who r lvl 222

OpenStudy (anonymous):

@rhondasommer

RhondaSommer (rhondasommer):

#YouCanOnlyGoToLevel99

OpenStudy (anonymous):

lol lies

OpenStudy (anonymous):

ask a real moderater he will tell u dif

RhondaSommer (rhondasommer):

@tkhunny can you help me out pleace

RhondaSommer (rhondasommer):

please*

OpenStudy (anonymous):

brb ima show u

OpenStudy (tkhunny):

This is why we don't post multiple questions on a single thread. Very confusing.